OpenStudy (anonymous):

which of the following is equivalent to x^5y^2/xy^2

OpenStudy (anonymous):

A.x^6y^5 B.x^5y c X^4y D. x^4

OpenStudy (anonymous):

either C or D... i think

OpenStudy (anonymous):

=x4y4

OpenStudy (anonymous):

its D

OpenStudy (anonymous):

Its C

OpenStudy (anonymous):

has anyone done this assignment ?

OpenStudy (mertsj):

x^5y^2/xy^2=x^(5-1)y^(2-2)= x^4y^0= x^4)(1)=x^4

OpenStudy (math&ing001):

You simplify by x*y^2 and you get x^4

OpenStudy (anonymous):

I said C

OpenStudy (anonymous):

it was d, thanks guys !(: